# Slice-aware Resource Allocation and Admission Control for Smart Factory Wireless Networks

### v2

@@ Abstract
The 5th generation (5G) and beyond network offers substantial promise as the ideal wireless technology to replace the existing inflexible wired connections in traditional factories of today. 5G network slicing allows for tailored allocation of resources to different network services, each with unique Quality of Service (QoS) requirements. This paper presents a novel solution for slice-aware radio resource allocation based on a convex optimisation control framework for applications in smart factory wireless networks. The proposed framework dynamically allocates minimum power and sub-channels to downlink mixed service type industrial users categorised into three slices: Capacity Limited (CL), Ultra Reliable Low Latency Communication (URLLC), and Time Sensitive (TS) slices. Given that the base station (BS) has limited transmission power, we enforce admission control by effectively relaxing the target rate constraints for current connections in the CL slice. This rate readjustment occurs whenever power consumption exceeds manageable levels. Simulation results show that our approach minimises power, allocates sub-channels to users, maintains slice isolation, and delivers QoS-specific communications to users in all the slices despite time-varying number of users and changing network conditions.
@@ I Introduction
Smart factory applications such as extendend reality (XR), robotics and automation, digital twins, and cyber-physical systems, leverage cutting-edge technologies to enhance manufacturing processes and improve overall factory efficiency [1][2]. One such technology is network slicing, a key feature of 5G wireless networks that utilise network functions virtualization (NFV) and software-defined networking (SDN) techniques to create logically isolated virtual networks (or ’slices’) on top of the same physical infrastructure [3]. Each slice functions as a distinct, dedicated end-to-end network and can be customised to meet the specific Quality-of-Service (QoS) requirements of different applications in terms of capacity, reliability, latency, and security. For example, one slice could be optimised for industrial IoT devices with low power and bandwidth needs, while another could support high-speed, low-latency applications like virtual reality and autonomous guided vehicles (AGVs)[4][5].
In an industrial setting, the complexity of managing shared radio resources is heightened by factors such as mobility of terminals, interference, varying propagation conditions, and diverse traffic types. These elements introduce uncertainty that must be carefully considered when allocating resources to the different slices, tailored for various smart factory applications. Ensuring that each slice meets its specific QoS requirements while maintaining strict isolation to prevent issues in one slice from impacting others presents a significant challenge. This is especially true given the demands for sub-millisecond latency and high reliability in the dynamic and interference-rich factory floor. Furthermore, with the constraints posed by limited radio resources [6] and the potential for high device density in smart factories [7], it is essential to deploy effective admission control mechanisms. These mechanisms are vital for managing network congestion and maintaining inter-slice priorities, thereby adding an extra layer of complexity to network management.
@@
  • Dynamically allocate DL radio resources such as power and subchannels to mixed traffic type users while prioritizing their unique QoS requirements. We model this problem by minimizing transmit power.
  • Significantly reduce the complexity of the optimisation problem formulation by translating all QoS constraints to target rate constraints.
@@ II System model
We classify users into three slice types depending on their specific applications and the type of traffic they generate:
@@
    Ultra-reliable low-latency communication (URLLC) slices. For applications such as augmented reality, real-time control of robots and machinery, AGVs, and real-time surveillance/security systems with stringent requirements on latency, jitter and reliability. This is typically asynchronous traffic characterized by a target outage delay given by $\Pr\left(D_{i}>D_{i}^{max}\right)<1-\gamma_{i}$ and short packet lengths, where $\gamma_{i}$ and $D_{i}^{max}$ are the target reliability and the maximum allowable delay respectively, for a user $i$ in a URLLC slice.
@@
    Time-sensitive (TS) slices. These slices cater to applications that demand immediate responses, such as machine vision systems or predictive maintenance. They are designed to handle periodically generated isochronous short packets that must be served upon arrival at the queue, with a target maximum bit error rate (BER). This is typically isochronous traffic with periodic arrivals at times that are multiples of the scheduling period $T_{s}$.
@@
The DL transmission rate from the BS to user $i$ in a Gaussian channel is upper bounded by the Shannon formula:
@@
where $B$ is the subchannel bandwidth and $h_{ij}(t)$ is the Rayleigh fading channel gain of the transmission from BS to user $i$ on subchannel $j$ at time slot $t$. $p_{ij}(t)$ is the DL transmission power of the BS on subchannel $j$ to user $i$ at slot $t$. $x_{ij}(t)$ is the subchannel allocation indicator with $x_{ij}(t)=1$ when subchannel $j$ is allocated to user $i$ at time slot $t$, else $x_{ij}(t)=0$. Finally, $\beta_{i}$ is a factor required to get a tight bound on the BER for an M-QAM modulation associated to user i [13]. Based on this model, the goal is to minimize the average power while maintaining slice isolation, reliability, latency and rates, a problem that we formally pose as:
@@
$r_{i}(t)=L_{i}/T_{s},\ \forall i\in\mathcal{N}_{s},\ \forall s\in\mathcal{S}_{T}$
@@
where $S_{D}$, $S_{C}$ and $S_{T}$ are the set of URLLC, CL and TS slices respectively. From the last constraint, we assume that each subchannel is allocated to at most one user and the target transmission rate for user $i$ is given by $r_{o,i}$. The objective function in (2) is the long-term average power spent by the BS. The first constraint is a reliability condition ensuring that delay is less than $D_{i}^{\text{max}}$ with a reliability of at least $1-\gamma_{i}$, thereby constituting the primary constraint for ensuring URLLC. The rate constraint is addressed implicitly using (1) and explicitly in the second and third constraints. Using this formulation, the instantaneous transmitted power on the j-th subchannel is $p_{j}(t)=\sum_{i=1}^{N}p_{ij}(t)$.
The problem in (2) is a stochastic mixed-integer linear programming optimization problem, which is NP-hard[9]. Moreover, as the number of users in the system varies over time, it is challenging to maintain slice isolation following a sudden change in the number of users. Thus, to simplify the problem, we convert the target outage delay constraint in (2) to a target rate constraint $r_{o,i}$. To that end, we find the probability with which the delay of user $i$ exceeds a threshold $D_{i}^{max}$, i.e., $\Pr\left(D_{i}>D_{i}^{max}\right)$. Assuming an M/M/1 queue, it can be proved that:
$\Pr(D_{i}>D_{i}^{max})=e^{-(r_{i}(t)-a_{i}(t))D_{i}^{max}}$
where $a_{i}(t)=\zeta_{i}(t)\beta_{i}(t)$ is the instantaneous arrival rate in bits per second (bps) for user $i$ which captures the instantaneous packet size $\zeta_{i}$ and the instantaneous packet arrival rate $\beta_{i}$ of user $i$ at time-slot $t$. It follows then that:
$e^{-(r_{i}(t)-a_{i}(t))D_{i}^{max}}<1-\gamma_{i}$
Finally, the rate constraint associated to a given outage delay of user $i$ at an instant of time can be written as:
@@
If we consider that the $i$-th user traffic has jitter constraints $J_{i}$, we can use the expression of the standard deviation of the exponential distribution assumed for $D$:
$std\{D\}=\frac{1}{r_{o,i}(\tau)-a_{s}(\tau)}.$ (4)
Therefore, a joint delay and jitter bound for a URLLC user yields the following required rate:
$r_{o,i}(t)=a_{i}(t)+\max\left(J^{-1}_{i},\frac{-\ln(1-\gamma_{i})}{D_{i}^{max}}\right).$ (5)
Also, knowing the packet size $L_{i}$ and the scheduling period $T_{s}$, the transmission rate of a user $i$ in a TS slice is given by:
$r_{o,i}=L_{i}/T_{s}.$ (6)
By imposing a rate constraint at each time instant, the average power optimization can be decoupled into instantaneous power optimization. We then recast the problem in (2) as:
@@ III-A Power
Problem (7) is convex if we relax $x_{ij}$ to take values in the interval $[0,1]$, so we rewrite the problem as $\min_{\mathbf{p},\mathbf{x}}\>\mathcal{L}$, where:
@@
By deriving $\mathcal{L}_{i}$ with respect to $p_{ij}$ we get the optimum power allocation for the $i$-th user:
@@
III-B Resource Blocks
To obtain the optimum variables $x_{ij}$ we plug (9) back into $\mathcal{L}_{i}(\mathbf{x},\mathbf{p};\boldsymbol{\lambda},\boldsymbol{\mu})$:
@@ III-B Resource Blocks
where
$\mu_{ij}(\lambda_{i})=\lambda_{i}B\log_{2}\left(1+\left[\frac{\lambda_{i}B}{\ln 2}-\frac{\sigma^{2}}{h_{ij}}\right]^{+}\frac{h_{ij}}{\sigma^{2}}\right)-\left[\frac{\lambda_{i}B}{\ln 2}-\frac{\sigma^{2}}{h_{ij}}\right]^{+}$
@@
Finally, we get $\mathcal{L}_{i}(\boldsymbol{\lambda},\boldsymbol{\mu})=\lambda_{i}r_{o,i}-\sum_{j=1}^{K}\left(\mu_{ij}(\lambda_{i})-\mu_{j}\right)^{+}$, and from (8) the dual function that must be maximized is:
@@
It can be proved (see appendix I) that $\mu_{j}^{*}=\max_{i}\mu_{ij}(\lambda_{i})$ minimizes the right term in (12), and hence:
@@
Note that the first condition in (11) will never occur. Additionally, the third condition states that any value for $x_{ij}$ is valid for the selected user in the j-th subchannel. Maximizing this expression requires a multi-dimensional search over $\lambda_{1},\dots,\lambda_{N}$. Since (13) is concave (the dual function is the point-wise infimum of a family of affine functions of $\lambda_{i}$) in the maximization we can apply subgradient principles, like the ellipsoid method [14], which has the advantage of the fast convergence and no tuning of a learning rate. The subgradient of (13) reads:
@@
A proper initialization of matrix $\mathbf{D}$ in algorithm 1 requires that the solution for $\boldsymbol{\lambda}$ is included in the ellipsoid defined by $\mathbf{D}$. If we can upperbound the dual variable as $\lambda_{i}^{*}\leq\lambda^{u}_{i}$, then a reasonable initial value for the matrix is $\mathbf{D}=\sum_{i}{\lambda_{i}^{u}}^{2}\mathbf{I}$. By recognising that the dual problem is concave and non-linear, equating (14) to zero is a sufficient condition to obtain $\lambda_{i}^{*}$:
@@
Note that the computation of $\lambda_{i}^{*}$ requires knowing $\mathcal{J}_{i}$, so we cannot obtain $\boldsymbol{\lambda}$ in closed form, but we can get an upperbound of $\lambda_{i}^{*}$ by applying the bound between arithmetic and geometric means, and also take $|\mathcal{J}_{i}|=1$:
@@
This bound provides an initial matrix for any scenario of rates, number of users and channel gains.
@@ III-C Guaranteed Slice Isolation
In (7), each user has a rate constraint associated with some QoS. If all constraints are met, the resource allocation framework isolates the service of one user from the rest. To this end, we model (7) as a control system as shown in Fig.1. The controller is designed to adjust the rate allocation among users in a network slice so that the total transmission rate adheres to a set limit. This approach guarantees slice isolation for CL slices, ensuring that each CL slice operates within its capacity limits without affecting others. The controller generates control signal $\boldsymbol{u}$ such that $\boldsymbol{u}=[u_{1}\;u_{2}\;\ldots\;u_{Ns}]^{T}$ is the input parameter vector that controls the solution of the optimisation problem. Note that $u_{i}(t)=r_{o,i}(t)$, i.e., the target rate at timeslot $t$ for all users $i$ in a CL. Thus, we assume that an external agent, based on a control system, recommends an instantaneous value for $r_{o,i}$ according to the observed rate of the delivered traffic.
Figure 1: Optimisation control system.
The control system equation reads;
$u_{s}(t)=u_{s}(t-1)-k\left(\sum_{i=1}^{N_{s}(t)}r_{i}(t)-C_{s}\right).$ (17)
At each time slot, the controller first calculates the deviation of the actual sum rate from the desired capacity limit, given by $\sum_{i=1}^{N_{s}(t)}r_{i}(t)-C_{s}$. This represents the error in the total rate allocation of the current slice. The controller then adjusts the target rate $u_{s}(t)$ based on this error scaled by $k<1/N_{s}(t)$, and the direction of the adjustment depends on whether the actual rate is above or below the capacity limit. This adjusted rate $u_{s}(t)$ becomes the new target rate $r_{o,i}$ for the next optimization at time $t+1$, helping to gradually steer the total rate of the CL slice towards $C_{s}$.

### v3

@@ Abstract
The 5th generation (5G) and beyond network offers substantial promise as the ideal wireless technology to replace the existing inflexible wired connections in traditional factories of today. 5G network slicing allows for tailored allocation of resources to different network services, each with unique Quality of Service (QoS) requirements. This paper presents a novel solution for slice-aware radio resource allocation based on a convex optimisation and control framework for applications in smart factory wireless networks. The proposed framework dynamically allocates minimum power and sub-channels to downlink mixed service type industrial users categorised into three slices: Capacity Limited (CL), Ultra Reliable Low Latency Communication (URLLC), and Time Sensitive (TS) slices. Given that the base station (BS) has limited transmission power, we enforce admission control by effectively relaxing the target rate constraints for current connections in the CL slice. This rate readjustment occurs whenever power consumption exceeds manageable levels. Simulation results show that our approach minimises power, allocates sub-channels to users, maintains slice isolation, and delivers QoS-specific communications to users in all the slices despite time-varying number of users and changing network conditions.
@@ I Introduction
Advanced technologies such as extended reality (XR), robotics and automation, digital twins, and cyber-physical systems, are pivotal in the development of smart factories, and can significantly boost the efficiency of manufacturing processes[1][2]. One of such innovative technologies is network slicing, a key feature of 5G wireless networks that utilise network functions virtualization (NFV) and software-defined networking (SDN) techniques to create logically isolated virtual networks (or ’slices’) on top of the same physical infrastructure [3]. Each slice functions as a distinct, dedicated end-to-end network and can be customised to meet the specific Quality-of-Service (QoS) requirements of different applications in terms of capacity, reliability, latency, and security. For example, one slice could be optimised for industrial IoT devices with low power and bandwidth needs, while another could support high-speed, low-latency applications like virtual reality and autonomous guided vehicle (AGV)[4].
In an industrial setting, the complexity of managing shared radio resources is heightened by factors such as mobility of terminals, interference, varying propagation conditions, and diverse traffic types. These elements introduce uncertainty that must be carefully considered when allocating resources to the different slices tailored for various smart factory applications. Ensuring that each slice meets its specific QoS requirements while maintaining strict isolation to prevent issues in one slice from impacting others presents a significant challenge. This is especially true given the demands for sub-millisecond latency and high reliability in the dynamic and interference-rich factory floor. Furthermore, with the constraints posed by limited radio resources [5] and the potential for high device density in smart factories [6], it is essential to deploy effective admission control mechanisms. These mechanisms are vital for managing network congestion and maintaining inter-slice priorities, thereby adding an extra layer of complexity to network resource management.
@@
  • Dynamically allocate DL radio resources such as power and subchannels to mixed traffic type users while prioritizing their unique QoS requirements. We model this problem by minimizing transmit power under rates constraints.
  • Significantly reduce the complexity of the optimisation problem formulation by translating all QoS constraints to target rate constraints. This translation allows for a efficient use of convex optimization algorithms, unlike in [8], [9].
@@ II System model
We classify users into three slice types depending on their specific application and the type of traffic they generate. Many smart factory use cases can be accommodated in these slice categories[12][13]:
@@
    Ultra-reliable low-latency communication (URLLC) slices. For applications such as augmented reality, real-time control of robots and machinery, AGV, and real-time surveillance/security systems with stringent requirements on latency, jitter and reliability. This is typically asynchronous traffic characterized by a target outage delay given by $\Pr\left(D_{i}>D_{i}^{max}\right)<1-\gamma_{i}$ and short packet lengths, where $\gamma_{i}$ and $D_{i}^{max}$ are the target reliability and the maximum allowable delay respectively, for a user $i$ in a URLLC slice.
@@
    Time-sensitive (TS) slices. These slices cater to applications that demand immediate responses, such as machine vision systems or predictive maintenance. They are designed to handle periodically generated isochronous short packets that must be served upon arrival at the queue, with a target maximum bit error rate (BER). Typically, these packets arrives at the queue at periodic intervals that are multiples of the scheduling period $T_{sp}$, with a fixed packet size $L$.
@@
The DL transmission rate from the BS to user $i$ in a Gaussian channel is upper bounded by the Shannon formula
@@
where $B$ is the subchannel bandwidth and $h_{ij}(t)$ is the Rayleigh fading channel gain of the transmission from BS to user $i$ on subchannel $j$ at time slot $t$. $\sigma^{2}$ is the noise power. Note that the interference levels across subchannels is captured in $\sigma^{2}$. $p_{ij}(t)/x_{ij}(t)$ is the DL transmission power on subchannel $j$ to user $i$ at slot $t$. $x_{ij}(t)$ is the subchannel allocation indicator with $x_{ij}(t)=1$ when subchannel $j$ is allocated to user $i$ at time slot $t$, else $x_{ij}(t)=0$. The presence of $x_{ij}(t)$ in the denominator of (1) guarantees the joint concavity of $r_{i}$ with respect to $x_{ij}(t)$ and $p_{ij}(t)$ [15]. Finally, $\beta_{i}$ is a factor required to get a tight bound on the uncoded BER for an M-QAM modulation associated to user i [14]. Based on this model, the goal is to minimize the average power while maintaining slice isolation, reliability, latency and rates, a problem formally posed as
@@
$r_{i}(t)=L/T_{sp},\ \forall i\in\mathcal{N}_{s},\ \forall s\in\mathcal{S}_{T}$
@@
where $S_{D}$, $S_{C}$ and $S_{T}$ are the set of URLLC, CL and TS slices respectively. The objective function in (2) is the long-term average power spent by the BS. The first constraint is a reliability condition ensuring that delay is less than $D_{i}^{\text{max}}$ with a reliability of at least $1-\gamma_{i}$, thereby constituting the primary constraint for ensuring URLLC. The rate constraint is addressed implicitly using (1) and explicitly in the second and third constraints. And in the last constraint, we assume that each subchannel is allocated to at most one user. Using this formulation, the instantaneous transmitted power on the j-th subchannel is $p_{j}(t)=\sum_{i=1}^{N}p_{ij}(t)$.
The problem in (2) is a stochastic mixed-integer linear programming optimization problem, which is NP-hard[8]. Moreover, as the number of users in the system varies over time, it is challenging to maintain slice isolation following a sudden change in the number of users. Thus, to simplify the problem, we convert the target outage delay constraint in (2) to a target rate constraint $r_{o,i}$. To that end, we find the probability with which the delay of URLLC user $i$ exceeds a threshold $D_{i}^{max}$, i.e., $\Pr\left(D_{i}>D_{i}^{max}\right)$. Assuming an M/M/1 queue, it can be proved that
$\Pr(D_{i}>D_{i}^{max})$ $=e^{-(r_{i}(t)-a_{i}(t))D_{i}^{max}},$ $\forall i\in\mathcal{N}_{s},\>\forall s\in\mathcal{S}_{D},$
where $a_{i}(t)=\zeta_{i}(t)\xi_{i}(t)$ is the instantaneous arrival rate in bits per second (bps) for user $i$ which captures the instantaneous packet size $\zeta_{i}$ and the instantaneous packet inter-arrival rate $\xi_{i}$ of URLLC user $i$ at time slot $t$. It follows then that
$e^{-(r_{i}(t)-a_{i}(t))D_{i}^{max}}<1-\gamma_{i}.$
Finally, the rate constraint associated to a given outage delay of user $i$ at an instant of time can be written as
@@
If we consider that the $i$-th user traffic has jitter constraints $J_{i}$, we can use the expression of the standard deviation of the exponential distribution assumed for $D$
$std\{D\}=\frac{1}{r_{o,i}(t)-a_{s}(t)}.$ (4)
Therefore, a joint delay and jitter bound for a URLLC user yields the following required rate
$r_{o,i}(t)$ $=a_{i}(t)+\max\left(J^{-1}_{i},\frac{-\ln(1-\gamma_{i})}{D_{i}^{max}}\right),$ (5)
$\quad\forall i\in\mathcal{N}_{s},\quad\forall s\in\mathcal{S}_{D},\quad\forall t.$
Also, knowing the packet size $L$ and the scheduling period $T_{sp}$, the transmission rate of a user $i$ in a TS slice is given by
$r_{o,i}=L/T_{sp},\quad\forall i\in\mathcal{N}_{s},\quad\forall s\in\mathcal{S}_{T}.$ (6)
By imposing a rate constraint at each time instant, the average power optimization can be decoupled into instantaneous power optimization. And the target transmission rate for any user $i$ is given by $r_{o,i}$. We then recast the problem in (2) as
@@ III-A Power
Problem (7) is convex and feasible if we relax $x_{ij}$ to take values in the interval $[0,1]$, so we rewrite the problem as $\min_{\mathbf{p},\mathbf{x}}\>\mathcal{L}$, where
@@
and $\lambda_{i}$ and $\mu_{j}$ are dual variables associated with the rate and subchannel allocation constraints, respectively. By deriving $\mathcal{L}_{i}$ with respect to $p_{ij}$ we get the optimum power allocation for the $i$-th user:
@@
III-B Subchannels
To obtain the optimum variables $x_{ij}$ we plug (9) back into $\mathcal{L}_{i}(\mathbf{x},\mathbf{p};\boldsymbol{\lambda},\boldsymbol{\mu})$
@@ III-B Subchannels
where we define the auxiliary variable
$\mu_{ij}(\lambda_{i})$ $=\lambda_{i}B\log_{2}\left(1+\left[\frac{\lambda_{i}B}{\ln 2}-\frac{\sigma^{2}}{h_{ij}}\right]^{+}\frac{h_{ij}}{\sigma^{2}}\right)$
$\quad-\left[\frac{\lambda_{i}B}{\ln 2}-\frac{\sigma^{2}}{h_{ij}}\right]^{+}$
@@
Finally, we get $\mathcal{L}_{i}(\boldsymbol{\lambda},\boldsymbol{\mu})=\lambda_{i}r_{o,i}-\sum_{j=1}^{K}\left(\mu_{ij}(\lambda_{i})-\mu_{j}\right)^{+}$, and from (8) the dual function that must be maximized is
@@
It can be proved (see appendix I) that $\mu_{j}^{*}=\max_{i}\mu_{ij}(\lambda_{i})$ minimizes the right term in (12), and hence
@@
Note that the first condition in (11) will never occur. Additionally, the third condition states that any value for $x_{ij}$ is valid for the selected user in the j-th subchannel. Maximizing this expression requires a multi-dimensional search over $\lambda_{1},\dots,\lambda_{N}$. Since (13) is concave (the dual function is the point-wise infimum of a family of affine functions of $\lambda_{i}$) in the maximization we can apply subgradient principles in the ellipsoid method [15], which has the advantage of the fast convergence and no tuning of a learning rate. The subgradient of (13) reads
@@
A proper initialization of algorithm 1 requires that the solution for $\boldsymbol{\lambda}$ is included in the ellipsoid defined by $\mathbf{D}$. If we can upper bound the dual variable as $\lambda_{i}^{*}\leq\lambda^{u}_{i}$, then a reasonable initial value is $\mathbf{D}=\sum_{i}{\lambda_{i}^{u}}^{2}\mathbf{I}$. By recognising that the dual problem is concave and non-linear, equating (14) to zero is a necessary and sufficient condition to obtain $\lambda_{i}^{*}$
@@
Note however that we cannot obtain $\boldsymbol{\lambda}^{*}$ in closed form since it requires $\mathcal{J}_{i}$, but we can upper bound it by applying the inequality between the arithmetic and the geometric means, extend the summation to all subchannels and take $|\mathcal{J}_{i}|=1$
@@
This bound provides an initial $\mathbf{D}$ given a scenario of transmission rates, number of subchannels and channel gains.
@@ III-C Guaranteed Slice Isolation
In (7), each user has a rate constraint associated with some QoS. If all constraints are met, the resource allocation framework isolates the service of one user from the rest. To this end, we model (7) as a control system as shown in Fig.1. The controller is designed to adjust the rate allocation among users in a network slice so that the total transmission rate adheres to a set limit. This approach guarantees slice isolation for CL slices, ensuring that each CL slice operates within its capacity limits without affecting others. The controller generates control signal $\boldsymbol{r_{o}}$ such that $\boldsymbol{r_{o}}=[r_{o,1}\;r_{o,2}\;\ldots\;r_{o,Ns}]^{T}$ is the input parameter vector that controls the solution of the optimisation problem. Where $r_{o,i}(t)$, is the target rate at time slot $t$ for a user $i$ in a CL slice. Thus, we assume that an external agent, based on a control system, recommends an instantaneous value for $r_{o,i}$ according to the observed rate of the delivered traffic.
Figure 1: Optimisation control loop for the CL slice
The control system equation reads
$r_{o,i}(t)=r_{o,i}(t-1)-k\left(\sum_{i=1}^{N_{s}(t)}r_{i}(t-1)-C_{s}\right).$ (17)
At each time slot, the controller first calculates the deviation of the actual sum rate from the desired capacity limit, given by $\sum_{i=1}^{N_{s}(t)}r_{i}(t-1)-C_{s}$. This represents the error in the total rate allocation at the current time, $t-1$ for the given CL slice $s$. The controller then adjusts the target rate $\boldsymbol{r_{o}}(t-1)$ based on this error scaled by $k<1/N_{s}(t)$, and the direction of the adjustment depends on whether the actual rate is above or below the capacity limit. This adjusted rate $\boldsymbol{r_{o}}(t)$ becomes the new target rate for the next optimization at time $t$, helping to gradually steer the total rate of the CL slice towards $C_{s}$.